# Gromov–Witten/Pandharipande–Thomas correspondence via conifold transitions

### v2

@@ 1. Introduction
Following the strategy in [LY05, HL12, LLW18], we will fit $X$ and $Y$ into two degenerations whose special fibers have a component in common and apply the degeneration formulas [Li02, MNOP06b, LW15].
@@ 2.1.2. Relative theories
Let $D$ be a smooth (not necessarily connected) divisor on $M$. Relative GW and PT theories enumerate curves with specified tangency to the divisor $D$. To impose the boundary conditions along $D$, we need the notion of cohomology weighted partitions.
@@
which parametrizes $r$-pointed relative stable maps of (arithmetic) genus $g\in\mathbb{Z}$ and degree $\beta$ with possibly disconnected domain curves and relative multiplicities determined by $\accentset{\rightharpoonup}{\eta}_{1},\cdots,\accentset{\rightharpoonup}{\eta}_{k}$. An element in the moduli space is a map to the stack of expanded relative pairs. As usual, a relative stable map has nonzero degrees on every connected component of its domain. It carries a virtual fundamental class of (complex) dimension
@@
The following rationality statement here is parallel to the absolute case [MNOP06b, PP14, PP17].
@@ 2.2. GW/PT correspondence
The key to relate descendent GW and PT-invariants, which are very different in flavor, is the correspondence matrices found by Pandharipande and Pixton [PP14, PP17]. See also [Obe24, §1.4 & 5.4]. The matrices relating GW and DT-invariants were predicted in [MNOP06b, Conj. 4].
@@ 2.2.1. Absolute version
If $\widehat{\alpha}=(\widehat{\alpha}_{1})$, then $\tau_{[\alpha]}(\gamma)=\tau_{\alpha_{1}-1}(\gamma)$.
@@ 2.3. The degeneration formulas
For the degeneration formulas in symplectic geometry, see [LR01, IP04, FTZ20].
@@ 3.2. GW/PT correspondence
    Let $\beta\in\operatorname{NE}(X)$ be a nonzero class and $\gamma_{1},\cdots,\gamma_{r}\in H^{\ast}(\mathfrak{X})$ be even classes. Suppose if $\gamma_{i}\in H^{0}(\mathfrak{X})$, then $\alpha_{i}=1$. Then we have
@@
(a) By the string equation for GW-invariants ([HKK+03, §26.3]), if the descendent insertions involve $\tau_{0}(1)$, then the calculation can be reduced to one with fewer insertions, removing this $\tau_{0}(1)$. On the PT-invariants side, by the string equation for PT-invariants ([Mor22, Prop. 1.1]), if the descendent insertions involve $\tau_{0}(1)$, then the PT-invariant is zero. Thus we can assume $\gamma_{i}\in H^{>0}(\mathfrak{X})$. We only need to consider cohomology classes of degrees $2,4$ and $6$. We consider the degeneration (3.2). Since the map from $Q_{j}$ to $\mathfrak{X}$ factors through $\{p_{j}\}\hookrightarrow\mathfrak{X}$ [LY06, §1], the pullback of $\gamma_{i}\in H^{>0}(\mathfrak{X})$ to $Q_{j}$ is zero for all $i$ and $j$. Thus the degeneration formulas will simplify so that there are no descendent insertions over $Q_{j}$.
@@
According to the virtual dimensions of the moduli spaces of relative stable maps (2.1),
@@
Since $E_{j}\subseteq Q_{j}$ is a hyperplane section, $\beta_{j}\cdot E_{j}=0$ implies that $\beta_{j}=0$. Then, according to (3.6) and (3.7), there is a unique $\beta_{Y}\in H_{2}(Y,\mathbb{Z})$ such that $\phi^{!}\beta_{Y}=\tilde{\beta}$ and $\psi_{\ast}\beta_{Y}=\beta$. Thus, the degeneration formula simplifies to the desired form, using the diagram (3.7).
@@
Keeping the notation of Proposition 3.5 and its proof, we first prove the following correspondence (cf. Conjecture 2.14)
$(-q)^{-\mathsf{c}_{\tilde{\beta}}^{\tilde{Y}}/2}\operatorname{Z_{PT}}\left(\tilde{Y}/E;q\,\middle|\,\tau_{\alpha_{1}-1}(\gamma_{1})\cdots\tau_{\alpha_{r}-1}(\gamma_{r})\,\middle|\,\eta\right)_{\tilde{\beta}}$
$=$ $(-iu)^{\mathsf{c}_{\tilde{\beta}}^{\tilde{Y}}+\ell(\eta)-|\eta|}\operatorname{Z_{GW}^{\prime}}\left(\tilde{Y}/E;u\,\middle|\,\overline{\tau_{\alpha_{1}-1}(\gamma_{1})\cdots\tau_{\alpha_{r}-1}(\gamma_{r})}\,\middle|\,\eta\right)_{\tilde{\beta}}$ (3.12)
for all $\tilde{\beta}\in\operatorname{Im}(\phi^{!})\subseteq\operatorname{NE}(\tilde{Y})$ under the variable change $-q=e^{iu}$. Here $\gamma_{i}$, $i=1,\dots,r$, are pulled back to $\tilde{Y}$ via the composition $\tilde{Y}\to Y\to\mathfrak{X}$. Under the map $Y\to\mathfrak{X}$, the curve $C_{j}$ is mapped to $p_{j}\in\mathfrak{X}$. Thus, the pullbacks are restricted to zero on $C_{j}$. By Proposition 3.5 (b), we have
$\sum_{\beta_{Y}\in\operatorname{NE}(Y)}\operatorname{Z}_{\mathrm{PT}}\left(Y;q\,\middle|\,\tau_{\alpha_{1}-1}(\gamma_{1})\cdots\tau_{\alpha_{r}-1}(\gamma_{r})\right)_{\beta}v^{\beta_{Y}}$ (3.13)
@@
$\operatorname{PT}\left(Y\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)$ $\coloneqq\sum_{\beta_{Y}\in\operatorname{NE}(Y)}\operatorname{Z_{PT}}\left(Y;q\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)_{\beta_{Y}}v^{\beta_{Y}}\quad\mbox{and }$
$\operatorname{PT_{exc}}(Y)$ $\coloneqq\prod_{i=1}^{k}\left(\sum_{m_{i}\geqslant 0}\operatorname{Z_{PT}}(\widetilde{E}_{i}/E_{i};q)_{m_{i}[\widetilde{C}_{i}]}v^{m_{i}[\tilde{C_{i}}]}\right),$
and similarly for
$\operatorname{GW^{\prime}}\left(Y\,\middle|\,\overline{\prod_{j=1}^{r}\tau_{\alpha_{j}-1}(\gamma_{j})}\right)\mbox{ and }\operatorname{GW_{exc}^{\prime}}(Y).$
Since $\operatorname{Z}_{\mathrm{PT}}(\widetilde{E}_{i}/E_{i};q)_{0}=\operatorname{Z}_{\mathrm{GW}}^{\prime}(\widetilde{E}_{i}/E_{i};u)_{0}=1$, the generating series $\operatorname{PT_{exc}}(Y)$ and $\operatorname{GW_{exc}^{\prime}}(Y)$ are invertible in $\mathbb{Q}(\!(q)\!)[\![\operatorname{NE}(Y)]\!]$ and $\mathbb{Q}(\!(u)\!)[\![\operatorname{NE}(Y)]\!]$ respectively. Then we may rewrite (3.13) as
$\frac{\operatorname{PT}\left(Y\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)}{\operatorname{PT_{exc}}(Y)}=\sum_{\beta_{Y}^{\prime}\in\operatorname{NE}(Y)}\operatorname{Z_{PT}}\left({\widetilde{Y}}/E;q\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)_{\phi^{!}\beta_{Y}^{\prime}}v^{\beta_{Y}^{\prime}},$ (3.14)
@@
$\frac{(-q)^{-\mathsf{c}_{\beta}^{Y}/2}\operatorname{PT}\left(Y\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)}{\operatorname{PT_{exc}}(Y)}=\frac{(-iu)^{\mathsf{c}_{\beta}^{Y}}\operatorname{GW^{\prime}}\left(Y\,\middle|\,\overline{\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)}\right)}{\operatorname{GW_{exc}^{\prime}}(Y)},$
@@ 4.2. Double solids
Let $X\nearrow Y$ be the conifold transition as in Proposition 4.3. If $Y$ is Calabi–Yau, i.e., $\mathcal{L}=\mathcal{O}_{\mathbb{P}^{3}}(4)$, then $Y$ satisfies the $\mathrm{GW}/\mathrm{PT}$ correspondence by [PP17, Thm. 1] since it’s a smooth hypersurface in a product of projective spaces. For the others, we can also degenerate the weak Fano complete intersection threefolds $Y$ by a similar factoring argument in the proof of [PP17, Thm. 1]. Hence the $\mathrm{GW}/\mathrm{PT}$ correspondence holds for $Y$. Then Corollary 4.5 follows from Theorem 3.2. ∎
@@ 4.3. Concluding remark
Acknowledgments. We would like to thank Yukinobu Toda, Chin-Lung Wang, Baosen Wu and Zijun Zhou for helpful discussions. YL is supported by grants from the Fundamental Research Funds for the Central Universities and Applied Basic Research Programs of Science and Technology Commission Foundation of Shanghai Municipality (22JC1402700). SSW is supported by the National Science and Technology Council (NSTC) under grant number 111-2115-M-001-003-MY3 and thanks Institute of Mathematics at Academia Sinica for providing support and a stimulating environment.

### v3

@@ 1. Introduction
To prove the theorem, the key is Proposition 3.5. We briefly explain its content and the strategy. See §3.1 for the meanings of various symbols. Following the strategy in [HL12, LLW18, LR01a, LY05], we will degenerate $X$ to $\widetilde{Y}\cup(\sqcup_{i}Q_{i})$ and $Y$ to $\widetilde{Y}\cup(\sqcup_{i}\widetilde{E}_{i})$, and apply the degeneration formulas [Li02, LW15, MNOP06b]. It essentially follows from dimension counts that each $Q_{i}$ has no contribution to the invariants. On the other hand, $\widetilde{E}_{i}$ is the toric compactification of a local curve, for which the GW/PT correspondence is known. Thus, we can relate the invariants of $X$ to those of $Y$ via $\widetilde{Y}$.
@@ 2.1.2. Relative theories
Let $D$ be a smooth (not necessarily connected) divisor on $M$. Relative GW and PT theories enumerate curves with specified tangency to the divisor $D$. To impose the boundary conditions along $D$, we need the notion of cohomology weighted partitions (see also [MNOP06b, §3.1]).
@@
which parametrizes $r$-pointed relative stable maps of (arithmetic) genus $g\in\mathbb{Z}$ and degree $\beta$ with possibly disconnected domain curves and relative multiplicities determined by $\accentset{\rightharpoonup}{\eta}_{1},\cdots,\accentset{\rightharpoonup}{\eta}_{k}$. When the cohomology weighted partitions are empty, we omit them from the expression. An element in the moduli space is a map to the stack of expanded relative pairs. As usual, a relative stable map has nonzero degrees on every connected component of its domain. It carries a virtual fundamental class of (complex) dimension
@@
Again, if the cohomology weighted partitions $\eta_{j}$’s are empty, we will omit them from the expression. The following rationality statement here is parallel to the absolute case [MNOP06b, PP14, PP17].
@@ 2.2. GW/PT correspondence
Descendent GW and PT-invariants are very different in flavor. The key to relate them is the correspondence matrices found by Pandharipande and Pixton [PP14, PP17]. See also [Obe24, §1.4 & 5.4]. The matrices relating GW and DT-invariants were predicted in [MNOP06b, Conj. 4].
@@ 2.2.1. Absolute version
If $\widehat{\alpha}=(\widehat{\alpha}_{1})$, then $\tau_{[\widehat{\alpha}]}(\gamma)=\tau_{\widehat{\alpha}_{1}-1}(\gamma)$.
@@ 2.3. The degeneration formulas
For the degeneration formulas in symplectic geometry, see [LR01b, IP04, FTZ20].
@@ 3.2. GW/PT correspondence
    Let $\beta\in\operatorname{NE}(X)$ be a nonzero class and $\gamma_{1},\cdots,\gamma_{r}\in H^{\ast}(\mathfrak{X})$ be even classes. Suppose if $\gamma_{i}\in H^{0}(\mathfrak{X})$, then $\alpha_{i}=1$. Then we have [^3] In the four displayed equalities in this proposition, the curve classes do not intersect the corresponding divisors. The weighted cohomology partitions are empty. So, we have omitted the weighted partitions from the relative generating functions, as indicated in §2.1.2.
@@
(a) By the string equation for GW-invariants ([HKK+03, §26.3]), if the descendent insertions involve $\tau_{0}(1)$, then the calculation can be reduced to one with fewer insertions, removing this $\tau_{0}(1)$. On the PT-invariants side, by the string equation for PT-invariants ([Mor22, Prop. 1.1]), if the descendent insertions involve $\tau_{0}(1)$, then the PT-invariant is zero. Thus we can assume $\gamma_{i}\in H^{>0}(\mathfrak{X})$ and only need to consider cohomology classes of degrees $2$, $4$, and $6$. We consider the degeneration (3.2). Since the map from $Q_{j}$ to $\mathfrak{X}$ factors through $\{p_{j}\}\hookrightarrow\mathfrak{X}$ [LY06, §1], the pullback of $\gamma_{i}\in H^{>0}(\mathfrak{X})$ to $Q_{j}$ is zero for all $i$ and $j$. Thus the degeneration formulas will simplify so that there are no descendent insertions over $Q_{j}$.
@@
Since there are no descendant insertions over $Q_{j}$, according to the virtual dimensions of the moduli spaces of relative stable maps (2.1),
@@
In particular, all weighted cohomology partitions are empty. Since $E_{j}\subseteq Q_{j}$ is a hyperplane section, $\beta_{j}\cdot E_{j}=0$ implies that $\beta_{j}=0$. Then, according to (3.6) and (3.7), there is a unique $\beta_{Y}\in H_{2}(Y,\mathbb{Z})$ such that $\phi^{!}\beta_{Y}=\tilde{\beta}$ and $\psi_{\ast}\beta_{Y}=\beta$. Thus, the degeneration formula simplifies to the desired form, using the diagram (3.7).
@@
Keeping the notation of Proposition 3.5 and its proof, we first prove the following correspondence without weighted cohomology partitions (cf. Conjecture 2.14)
$(-q)^{-\mathsf{c}_{\tilde{\beta}}^{\tilde{Y}}/2}\operatorname{Z_{PT}}\left(\tilde{Y}/E;q\,\middle|\,\tau_{\alpha_{1}-1}(\gamma_{1})\cdots\tau_{\alpha_{r}-1}(\gamma_{r})\right)_{\tilde{\beta}}$
$=$ $(-iu)^{\mathsf{c}_{\tilde{\beta}}^{\tilde{Y}}}\operatorname{Z_{GW}^{\prime}}\left(\tilde{Y}/E;u\,\middle|\,\overline{\tau_{\alpha_{1}-1}(\gamma_{1})\cdots\tau_{\alpha_{r}-1}(\gamma_{r})}\right)_{\tilde{\beta}}$ (3.12)
for all $\tilde{\beta}\in\operatorname{Im}(\phi^{!})\subseteq\operatorname{NE}(\tilde{Y})$ under the variable change $-q=e^{iu}$. Here $\gamma_{i}$, $i=1,\dots,r$, are pulled back to $\tilde{Y}$ via the composition $\tilde{Y}\to Y\to\mathfrak{X}$. Under the map $Y\to\mathfrak{X}$, the curve $C_{j}$ is mapped to $p_{j}\in\mathfrak{X}$. Thus, the pullbacks are restricted to zero on $C_{j}$. By Proposition 3.5 (b) and $K_{Y}\cdot C_{j}=0$ for $1\leqslant j\leqslant k$, we have
$\sum_{\beta_{Y}\in\operatorname{NE}(Y)}(-q)^{-\mathsf{c}_{\beta_{Y}}/2}\operatorname{Z}_{\mathrm{PT}}\left(Y;q\,\middle|\,\tau_{\alpha_{1}-1}(\gamma_{1})\cdots\tau_{\alpha_{r}-1}(\gamma_{r})\right)_{\beta}v^{\beta_{Y}}$ (3.13)
@@
$\operatorname{PT}\left(Y\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)$ $\coloneqq\sum_{\beta_{Y}\in\operatorname{NE}(Y)}(-q)^{-\mathsf{c}_{\beta_{Y}}/2}\operatorname{Z_{PT}}\left(Y;q\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)_{\beta_{Y}}v^{\beta_{Y}},$
$\operatorname{PT_{exc}}(Y)$ $\coloneqq\prod_{i=1}^{k}\left(\sum_{m_{i}\geqslant 0}\operatorname{Z_{PT}}(\widetilde{E}_{i}/E_{i};q)_{m_{i}[\widetilde{C}_{i}]}v^{m_{i}[\tilde{C_{i}}]}\right).$
We define
$\operatorname{GW^{\prime}}\left(Y\,\middle|\,\overline{\prod_{j=1}^{r}\tau_{\alpha_{j}-1}(\gamma_{j})}\right)$ $\coloneqq\sum_{\beta_{Y}\in\operatorname{NE}(Y)}(-iu)^{\mathsf{c}_{\beta_{Y}}}\operatorname{Z_{GW}^{\prime}}\left(Y;q\,\middle|\,\overline{\prod_{j=1}^{r}\tau_{\alpha_{j}-1}(\gamma_{j})}\right)_{\beta_{Y}}v^{\beta_{Y}},$
and $\operatorname{GW_{exc}^{\prime}}(Y)$ similarly. Since $\operatorname{Z}_{\mathrm{PT}}(\widetilde{E}_{i}/E_{i};q)_{0}=\operatorname{Z}_{\mathrm{GW}}^{\prime}(\widetilde{E}_{i}/E_{i};u)_{0}=1$, the generating series $\operatorname{PT_{exc}}(Y)$ and $\operatorname{GW_{exc}^{\prime}}(Y)$ are invertible in $\mathbb{Q}(\!(q)\!)[\![\operatorname{NE}(Y)]\!]$ and $\mathbb{Q}(\!(u)\!)[\![\operatorname{NE}(Y)]\!]$ respectively. Then we may rewrite (3.13) as
$\frac{\operatorname{PT}\left(Y\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)}{\operatorname{PT_{exc}}(Y)}$ (3.14)
$=$ $\sum_{\beta_{Y}^{\prime}\in\operatorname{NE}(Y)}(-q)^{-\mathsf{c}_{\beta_{Y}^{\prime}}/2}\operatorname{Z_{PT}}\left({\widetilde{Y}}/E;q\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)_{\phi^{!}\beta_{Y}^{\prime}}v^{\beta_{Y}^{\prime}},$
@@
$\frac{\operatorname{PT}\left(Y\,\middle|\,\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)\right)}{\operatorname{PT_{exc}}(Y)}=\frac{\operatorname{GW^{\prime}}\left(Y\,\middle|\,\overline{\prod_{j=1}^{r}\tau_{\alpha_{j}-1}\left(\gamma_{j}\right)}\right)}{\operatorname{GW_{exc}^{\prime}}(Y)},$
@@ 4.2. Double solids
Let $X\nearrow Y$ be the conifold transition as in Proposition 4.3. If $Y$ is Calabi–Yau, i.e., $\mathcal{L}=\mathcal{O}_{\mathbb{P}^{3}}(4)$, then $Y$ satisfies the $\mathrm{GW}/\mathrm{PT}$ correspondence by [PP17, Thm. 1] since it is a smooth hypersurface in a product of projective spaces. For the others, we can also degenerate the weak Fano complete intersection threefolds $Y$ by a similar factoring argument in the proof of [PP17, Thm. 1]. Hence the $\mathrm{GW}/\mathrm{PT}$ correspondence holds for $Y$. Then Corollary 4.5 follows from Theorem 3.2. ∎
@@ 4.3. Concluding remark
Acknowledgments. We thank Yukinobu Toda, Chin-Lung Wang, Baosen Wu and Zijun Zhou for helpful discussions. We also thank the referees for helpful comments and suggestions on this article. YL is supported by grants from the Fundamental Research Funds for the Central Universities and Applied Basic Research Programs of Science and Technology Commission Foundation of Shanghai Municipality (22JC1402700). SSW is supported by the National Science and Technology Council (NSTC) under grant number 111-2115-M-A49-019-MY3 and thanks the Institute of Mathematics at Academia Sinica for providing support and a stimulating environment.